# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ VI Whole-body control
The WBC solves an instantaneous control problem (i.e., only for the current timestep). This control scheme employs Euler-Lagrangian dynamics to express equation (2) as
$M\boldsymbol{\ddot{q}}-J^{T}\boldsymbol{\lambda}-S^{T}\boldsymbol{\tau}=-C\boldsymbol{\dot{q}}-G$ (5)
@@
For a humanoid robot with a floating base, the rank of $S$ is smaller than the dimension of the generalized position $\boldsymbol{q}$. This implies that the humanoid system is underactuated; it requires physical contact with the environment to maintain balance and achieve mobility and manipulation tasks. The contact constraint is described using contact Jacobian $J_{c}$:
$J_{c}\boldsymbol{\dot{q}}=0\Rightarrow J_{c}\boldsymbol{\ddot{q}}+\dot{J}_{c}\boldsymbol{\dot{q}}=0$ (6)
These underactuated and contact-constrained dynamics (5), (6) represent the main ingredients in solving the WBC for humanoid robots.
@@ VI-A WBC Dynamic Tasks
where $t$ is the time. Dynamic tasks can be equality constraints ($e_{i}=0$), inequality constraints ($e_{i}\leq 0)$, or cost terms ($|e_{i}|^{2}$). The main idea of WBC is that the linear equation (7) is sufficient to universally describe a diverse set of locomotion and manipulation tasks.
@@
As shown in Fig. 9, to achieve a desired list of dynamic tasks, WBC approaches can be categorized based on closed-form or optimization-based approaches.
@@ VI-B WBC in Closed Form
An inverse dynamics controller is among the early works that address the WBC problem in closed form. In particular, it solves a single dynamic task: achieving the desired generalized acceleration $\boldsymbol{\ddot{q}}=\boldsymbol{\ddot{q}}^{d}$. The closed-form torque $\boldsymbol{\tau}$ can be solved from (5) if we can measure all constraint forces $\boldsymbol{\lambda}$. However, $\boldsymbol{\lambda}$ are usually unattainable due to the lack of sensing and estimation capability. To derive torque analytically, several methods [251, 252, 253] project the system dynamics (5) into a constraint-free manifold, establishing an equation between only $\boldsymbol{\ddot{q}}$ and $\boldsymbol{\tau}$. [251] project (5) into the actuated joint space, resulting in an equation of constrained dynamics independent of the constraint forces $\boldsymbol{\lambda}$[252] use orthogonal decomposition of the constraint Jacobian $J$ to project (5) into the nullspace of the constraints. Although different projection techniques are used, [253] show that these projections produce equivalent torque results. However, from the perspective of solving speed and numerical stability, orthogonal decomposition [252] is favored as it is faster and does not require the inversion of the inertia matrix.
@@ VI-C WBC through Optimization
In contrast, weighted QP addresses these issues via arbitrating dynamic tasks as soft constraints in cost functions, using weights to express their relative priority. Therefore, weighted QP can be regarded as a special case of hierarchical QP with only one hierarchy or vice-versa as documented in [262]. Such a setup enjoys the benefit of solving only a single optimization, which is faster than a hierarchical QP and can be further accelerated by exploiting sparsity and warm-start. However, tuning weight parameters can be burdensome for a large number of tasks, and can lead to instability [263]. Even with well-tuned parameters, the loss of strict task priorities means low-priority tasks can interfere with high-priority ones. Nevertheless, weighted QP is still widely applied in many robotics studies due to its easy setup and computation efficiency compared with hierarchical QP. For example, many of the weighted QP methods were designed during the DARPA Robotics Challenge [238, 264, 117, 237].
@@ VI-D1 Interaction as an External Wrench
The second strategy, known as pre-optimization [3, 266], involves two stages in sequence. First, it determines the optimal distribution of contact wrenches based on the desired rate of centroidal momentum derived from the balance stability of CD. The second stage computes the joint torques needed to realize the contact wrench using inverse dynamics of full-body dynamics. Note that, deriving the desired rate of centroidal momentum in the first stage is particularly challenging due to the non-holonomy [267] of angular momentum, i.e., the kinetic momentum of rotation is not directly related to the orientation of body links. As a result, the body orientation requires additional regulation (e.g., joint-level postural feedback  [3]) beyond simple feedback on angular momentum.
@@ VI-D2 Interaction as a Unified Robot-Object Model
Conclusion: The core of whole-body control lies in addressing an inverse dynamics problem to produce joint-level torque control. However, this problem is challenging due to the underactuation and contact-constrained nature of humanoid robots. Closed-form approaches such as inverse dynamics control and operational-space control are computationally efficient. Therefore, they have been traditionally prevalent. On the other hand, optimization-based strategies, particularly quadratic programs, are increasingly favored as they adapt more effectively to a wide range of task specifications and offer more reliable solutions. Undoubtedly, both lines of WBC research have significantly advanced the progress of humanoid robot control over the past two decades. In the near term, optimization-based WBC will continue to be a popular choice for low-level control to achieve high-level loco-manipulation tasks. We also see neural WBC [270, 271, 272] gaining popularity, as we will discuss in the following section. For further reading, we recommend the survey on optimization-based WBC for legged robots [7] and the chapter on closed-form WBC techniques for humanoid robots [11].
@@ VII Learning Loco-Manipulation Skills
This section reviews learning-based approaches that explore two main directions: (i) enhancing a specific skill in terms of agility, robustness, and safety, and (ii) broadening the overall skill set of robots, revolving around two pivotal goals: versatility and generalization. Versatility refers to the capability of a single framework or policy to master multiple skills, whereas generalization involves adapting existing skills to new, out-of-distribution tasks and environments.
@@ VII-A Skill Learning: Reinforcement Learning from Scratch
RL provides distinct benefits but also comes with its own challenges. Compared with model-based methods, numerous RL methods are model-free (see Sec. VII-E for model-based RL), eliminating the need for accurate dynamics. Furthermore, RL does not require demonstration data, making its training setup straightforward. However, it often requires meticulous design of reward functions to shape the policy’s behavior. In addition, deploying a learned policy on robot hardware often encounters a sim-to-real gap, a well-known issue induced by the inaccurate physical model used by the simulator. Policy learning from scratch often requires extensive and time-consuming interactions with the environment without a guarantee for task completion. For example, popular RL algorithms such as Proximal Policy Optimization (PPO) [278] and Soft Actor-Critic (SAC) [279] fail to achieve most humanoid loco-manipulation tasks [280], partly due to the complexity of these tasks, the sample inefficiency, and the sparse reward design.
@@
VII-A1 Improving Learning Efficiency
@@
Several approaches have been proposed to improve learning efficiency. Curriculum learning expedites training by allowing the policy to achieve simple tasks in the early stage of training and then progressively increasing the task difficulty and complexity [281]. Another approach is to promote exploration. Researchers use curiosity mechanisms, which encourage visiting unexplored states, to intrinsically motivate learning without explicit reward design [282]. This has been shown to overcome the sparse reward setting and achieve complex loco-manipulation behaviors such as door opening. [274] also incorporates curiosity-based rewards to learn versatile loco-manipulation skills without any motion priors. Lastly, substituting reward terms with constraints in a constrained RL framework can significantly simplify reward tuning while achieving state-of-the-art locomotion performance [283].
VII-A2 Addressing Sim-to-real Gap
@@
Sim-to-real is another formidable challenge in RL. Nevertheless, RL has been successfully applied to various areas of robotics, with notable success in quadrupeds [284], where the sim-to-real gap has been consistently overcome. How can we apply the lessons from quadrupeds to humanoids for optimal loco-manipulation while avoiding known pitfalls? The success story of quadrupeds hinges on new investments in infrastructure for affordable hardware and highly parallel physics engines, spearheaded by key players in the robotics community. It is also important to note that quadrupeds benefit from an inherently stable dynamic system similar to manipulators while operating in less complex environments compared to typical loco-manipulation tasks.
@@
In contrast, humanoid loco-manipulation faces steeper sim-to-real challenges. Humanoid robots possess higher DoFs and unstable dynamics, where the center of mass constantly moves out of the support polygon. Therefore, learning whole-body balance control is sensitive to parameters in physics simulation, underlining the sim-to-real gap due to dynamics differences between the virtual and real worlds. In addition, humanoids are expected to perform human-level manipulation tasks where the differences in observation space and the complex environment aggravate the sim-to-real-gap due to appearances.
@@
To address the sim-to-real challenge, a diverse set of mainstream approaches have been explored for humanoid robots. Domain Randomization (DR) is among the most widely adopted approaches. It varies the properties of a robot model, such as mass, friction, and actuator dynamics, to train a generalized policy robust in the real world. Many humanoid works [275, 39] achieve sim-to-real transfer through DR. While DR is straightforward to set up, policy training is sensitive to the parameter randomization range, inducing laborious tuning: a larger range is challenging for the policy to fit all physical parameters (i.e., fail to learn), and a smaller range does not cover the full spectrum of parameters that can occur in the real physical world (i.e., fail to transfer).
@@
In addition to adopting a diverse set of parameters in DR, System Identification (SI) is another popular approach to enhancing model fidelity by approximating the system’s input-output behavior from real-world data. Real-to-sim techniques use optimization [285] or search [286] to obtain simulation parameters that can best explain the real robot trajectories collected from policy execution. However, it is challenging to collect high-quality, real-world data covering the full space of states and actions, particularly for versatile, safety-critical humanoid tasks.
@@
Different from SI, which uses real-world data to obtain an accurate model, Domain Adaptation (DA) uses real-world data directly to fine-tune a simulator-trained policy. In DA, the parameter distribution in simulation is defined as the source domain, and parameter distribution in the real world is defined as the target domain. The fine-tuning effectively transfers the policy from the source domain to the target domain. For example, Sim-to-Lab-to-Real [287] develops a two-stage transfer: pre-training in simulation and fine-tuning in the real world. Although only limited hardware data is needed for fine-tuning, safety is still a major concern. Safety filters are often deployed to prevent dangerous movements when collecting real-world data [288].
Despite these efforts to address the sim-to-real gap, a systematic solution remains elusive, as the aforementioned approaches are often ad-hoc and case-specific. Against this backdrop, advancing in humanoid hardware deployment and developing physics engines that facilitate real-to-sim construction and sim-to-real transfer will be crucial. Simultaneously, we must rethink the role of RL to utilize its strengths and avoid its weaknesses for humanoid loco-manipulation. An in-depth discussion on how to integrate RL with modern IL is presented in Sec. VII-B and Sec. VII-C.
@@
Conclusion: RL provides an effective way to learn novel behaviors for humanoid loco-manipulation. However, in practice, the success of RL often relies on informative representations for both observation and action, extensive reward engineering, curriculum learning design, and a vast amount of trial-and-error experiences to estimate gradients. Consequently, using RL to train robots is almost never practical in the real world, at least for the current stage of development. Therefore, RL policies are predominantly trained in simulations. This makes the sim-to-real gap the Achilles’ heel of RL, significantly dampening its initial promise. Compared with quadruped robots, the sim-to-real gap is particularly challenging for humanoid robots with high DoFs executing complex loco-manipulation tasks. This is why IL, leveraging limited but in-domain real-world data, has gained popularity over pure RL without demonstration data. For further reading in RL, we recommend the survey on learning-based legged locomotion [277] and the practical lessons for training robotic RL agent [289].
TABLE III: Skill Learning Methods Based On Data Source

### v2

@@ VI Whole-body control
A WBC is different from an MPC in two aspects. (i) The MPC often solves a receding horizon problem, whereas the WBC solves an instantaneous control problem (i.e., only for the current timestep). (ii) The model adopted by MPC may vary based on the user choice, as discussed in Sec. V, but WBC usually employs full-order Euler-Lagrangian dynamics that express equation (2) as
$M\boldsymbol{\ddot{q}}-J^{T}\boldsymbol{\lambda}-S^{T}\boldsymbol{\tau}=-C\boldsymbol{\dot{q}}-G,$ (5)
@@
In a humanoid robot with a floating base, the rank of $S$ is smaller than the dimension of the generalized position $\boldsymbol{q}$, meaning the system is underactuated. This requires physical contact with the environment to achieve balance, mobility, and manipulation. The contact constraint is described using contact Jacobian $J_{c}$:
$J_{c}\boldsymbol{\dot{q}}=0\Rightarrow J_{c}\boldsymbol{\ddot{q}}+\dot{J}_{c}\boldsymbol{\dot{q}}=0.$ (6)
These underactuated and contact-constrained dynamics (5), (6) represent the main components for solving the WBC for humanoid robots.
@@ VI-A WBC Dynamic Tasks
where $t$ is the time. Dynamic tasks can be equality constraints ($e_{i}=0$), inequality constraints ($e_{i}\leq 0)$, or cost terms ($|e_{i}|^{2}$). The main idea of WBC is that the linear equation (7) is sufficient to describe a universal set of locomotion and manipulation tasks.
@@
As shown in Fig. 9, WBC approaches can be categorized based on closed-form or optimization-based approaches to achieve a desired list of dynamic tasks.
@@ VI-B WBC in Closed Form
An inverse dynamics controller is among the early works that address the WBC problem in closed form. In particular, it solves a single dynamic task: achieving the desired generalized acceleration $\boldsymbol{\ddot{q}}=\boldsymbol{\ddot{q}}^{d}$. The closed-form torque $\boldsymbol{\tau}$ can be solved from (5) if we can measure all constraint forces $\boldsymbol{\lambda}$. However, $\boldsymbol{\lambda}$ is usually unattainable due to the lack of sensing and estimation capability. To derive torque analytically, several methods [247, 248, 249] project the system dynamics (5) into a constraint-free manifold, establishing an equation between only $\boldsymbol{\ddot{q}}$ and $\boldsymbol{\tau}$. [247] projects (5) into the actuated joint space, resulting in an equation of constrained dynamics independent of the constraint forces $\boldsymbol{\lambda}$[248] uses orthogonal decomposition of the constraint Jacobian $J$ to project (5) into the nullspace of the constraints. Although different projection techniques are used, [249] shows that these projections produce equivalent torque results. However, from the perspective of solving speed and numerical stability, orthogonal decomposition [248] is favored as it is faster and does not require the inversion of the inertia matrix.
@@ VI-C WBC through Optimization
In contrast to hierarchical QP, which solves tasks in a strict order, weighted QP resolves task conflicts by treating dynamic tasks as soft constraints within the cost functions, with weights indicating their relative priority. Therefore, weighted QP can be regarded as a special case of hierarchical QP with only one hierarchy or vice-versa as documented in [258]. Such a setup enjoys the benefit of solving only a single optimization, which is faster than a hierarchical QP and can be further accelerated by exploiting sparsity and warm-start. However, tuning weight parameters can be burdensome for a large number of tasks, and can lead to instability [259]. Even with well-tuned parameters, the loss of strict task priorities means low-priority tasks can interfere with high-priority ones. Nevertheless, weighted QP is still widely applied in many robotics studies due to its easy setup and computation efficiency compared with hierarchical QP. For example, many of the weighted QP methods were designed during the DARPA Robotics Challenge [234, 260, 130, 233].
@@ VI-D1 Interaction as an External Wrench
The second strategy, known as pre-optimization [3, 262], involves two stages in sequence. First, it determines the optimal distribution of contact wrenches based on the desired rate of centroidal momentum derived from the balance stability of CD. The second stage computes the joint torques needed to realize the contact wrenches using inverse dynamics of full-body dynamics. Note that, deriving the desired rate of centroidal momentum in the first stage is particularly challenging due to the non-holonomy [263] of angular momentum, i.e., the kinetic momentum of rotation is not directly related to the orientation of body links. As a result, the body orientation requires additional regulation (e.g., joint-level postural feedback  [3]) beyond simple feedback on angular momentum.
@@ VI-D2 Interaction as a Unified Robot-Object Model
Key Takeaways for Whole-Body Control: The core of whole-body control lies in addressing an inverse dynamics problem to produce joint-level torque control. However, this problem is challenging due to the underactuation and contact-constrained nature of humanoid robots. Closed-form approaches such as inverse dynamics control and operational-space control are computationally efficient. Therefore, they have been traditionally prevalent. On the other hand, optimization-based strategies, particularly quadratic programs, are increasingly favored as they adapt more effectively to a wide range of task specifications and offer more reliable solutions. Undoubtedly, both lines of WBC research have significantly advanced the progress of humanoid robot control over the past two decades. In the near term, optimization-based WBC will continue to be a popular choice for low-level control to achieve high-level loco-manipulation tasks. We also see neural WBC [266, 267, 268] gaining popularity, as we will discuss in the following section. For further reading, we recommend the survey on optimization-based WBC for legged robots [9] and the chapter on closed-form WBC techniques for humanoid robots [7].
VI-E Challenges in Numerical Optimization
Robotic planning and control techniques that are formalized as numerical optimization problems heavily rely on advances in applied discrete mathematics and optimization theory. These advancements address challenges such as nonconvexity, numerical robustness, and real-time resolution performance. However, a plateau may have been reached in the transfer of these techniques to the field of robotics at large – humanoids specifically – despite exploiting the unique properties of robotic models. These properties can enhance the efficiency of optimization problem formulation and its resolution by tailoring them to specific applications. However, as evidenced by the formulation OCP, including MPC (Sec. V) and WBC (Sec. VI), inherent physical uncertainties can disrupt closed-loop performance. Extending these formulations to loco-manipulation primarily involves (i) augmenting models to incorporate loco-manipulated counterparts and (ii) refining contact models formulations to account for various interactions (e.g., impact [269], rolling, deforming). However, these extensions risk overcomplicating the problem, potentially hindering effective formalization, even if the resulting formulations are sparse.
Contact-explicit optimization formulations [6] are generally preferred due to their faster convergence and simplified formulation. However, they still suffer from the curse of dimensionality, which limits the complexity of the considered problem, the length of the preview window, and the discretization granularity of the decision variables. While increasing the resolution of the discretization and the length of the preview horizon increases the quality and stability of the results, the corresponding decrease in solution speed can, in some cases, cause an overall reduction in system stability. However, as computational power increases, this lowers the significance of these challenges.
Importantly, however, the contact-explicit formulations have the significant limitation of requiring the user to determine the contact mode sequence for the problem, which generally limits the ability to generate complex motions. Alternatively, contact-implicit formulations introduce complementarity conditions to eliminate the strict dependence on the contact mode sequence [162, 270, 167]. However, contact complementarity conditions are nonsmooth, introducing severe computational challenges. Generally, this is tackled via regularizing the complementarity problem, e.g.[170], which approximates the constraint with a continuous affine function. Even with this linearized approximation, contact-implicit approaches struggle to scale to the high dimensionality of humanoid robots due to excessive computation and numerical difficulty. However, improved solver formulations, either as first-order solvers such as Alternating Direction Method of Multipliers (ADMM) [271] or specialized methods [168] for handling the complementarity conditions, have the potential to increase the tractability of real-time contact-implicit optimization.
Possibly more importantly, all these approaches, however, still have only local optimality guarantees. The level of nonconvexity or nonlinearity of the problem influences the solution quality, which often results in poor local minima or sub-optimal solutions. When the structure of the problem requires deviations from local candidate solutions, a feasible solution, even one that is only locally optimal, may never be found. This feasibility issue remains an open challenge for nonconvex nonlinear optimization.
Additionally, optimization approaches are almost always deterministic in nature, failing to capture the stochasticity in the state estimates and future contact events. This lack of consideration of uncertainty along the optimization horizon is critical, as it leads to limited robustness in translation to the real world. Addressing this uncertainty and lack of global optimality has led to the combination of search techniques with traditional trajectory optimization. For example, Model Predictive Path Integral (MPPI) [211] samples a variety of random control signals and their resulting state trajectories to determine the best action to take. Alternative contacts and objectives can also be sampled with contact-implicit approaches to help avoid local minima and find the globally optimal solution to accomplish a task [272]. Although both of these approaches heavily leverage computational parallelization for expediency, parallelizing the underlying optimization algorithms explicitly designed for trajectory optimization is gaining increasing prevalence, both on the CPU [273] and GPU [274, 205], as discussed in Sec. V. Despite this, the inclusion of the uncertainty into the optimization problem, whether through sampling [275], smoothing [276], or the introduction of stochasticity [277], is a promising approach for mitigating some of these challenges.
Despite the gains in these algorithms for considering the full system dynamics, the robustness of the mathematical solution when performing numerical optimization has been a concern due to the infeasibility of complex optimization problems, regardless of the computation speed improvement. As discussed in Sec. VI WBC, arbitrating techniques address infeasibility by (i) relaxing the hard constraint to the soft constraint by combining them in the cost with a weighted sum and (ii) prioritizing the constraints by achieving the important ones first. However, how to design a smart solver to automatically resolve this issue and provide numerical robustness is still an open question. In addition, weight-tuning in high-dimensional problems with complex objectives is nontrivial, highly task-dependent, and can lead to instability [259]. Researchers have made initial steps to apply auto-tuning techniques to streamline the tuning process in Optimal Control Problem (OCP) designs for humanoid robots [278, 279]. Until these are solved, assigning definitive but non-violating constraints and designing objective functions while maintaining global versatility still depend on expert knowledge.
@@ VII Learning Loco-Manipulation Skills
This section reviews learning-based approaches that explore two main directions: (i) enhancing a specific skill in terms of agility, robustness, and safety, and (ii) broadening the overall skill set of robots, revolving around two pivotal goals: versatility and generalization. Versatility refers to the capability of a single framework or policy to master multiple skills, whereas generalization involves adapting existing skills to new, out-of-distribution tasks and environments. Versatility can be viewed as a prerequisite for generalization. For a model to generalize across a wide range of tasks, it must possess a broad and versatile set of skills that can be applied in novel contexts. However, possessing versatile skills alone does not guarantee generalization—the model must also be capable of correctly prompting and adapting those skills to meet novel objectives. For instance, in the context of skill learning, a versatile policy can track multiple trajectories within the training dataset, while a generalizable policy should be able to track motions beyond those seen in the dataset, thereby demonstrating out-of-distribution capabilities.
@@ VII-A Skill Learning: Reinforcement Learning from Scratch
RL provides distinct benefits but also comes with its own challenges. Compared with model-based methods, numerous RL methods are model-free (see Sec. VII-E for model-based RL), eliminating the need for accurate dynamics. Furthermore, RL does not require demonstration data, making its training setup straightforward. However, it often requires meticulous design of reward functions to shape the policy’s behavior. In addition, deploying a learned policy on robot hardware often encounters a sim-to-real gap, a well-known issue induced by the inaccurate physical model used by the simulator. Policy learning from scratch often requires extensive and time-consuming interactions with the environment without a guarantee for task completion. For example, popular RL algorithms such as Proximal Policy Optimization (PPO) [284] and Soft Actor-Critic (SAC) [285] fail in most humanoid loco-manipulation tasks [286], partly due to the complexity of these tasks, the sample inefficiency, and the sparse reward design.
@@
VII-A1 Comparison Between RL Algorithms
PPO [284], and similarly other on-policy reinforcement learning algorithms such as A3C [287], is sample inefficient, but the policy update is fast and scalable. PPO is widely used in sim-to-real settings where simulation data is relatively cheap to obtain, e.g., in a parallelized simulation environment such as Isaac Gym [215]. SAC [285], and similarly, Twin Delayed Deep Deterministic (TD3) [288] and Deep Deterministic Policy Gradient (DDPG) [289], are sample-efficient off-policy algorithms, but slower to train due to higher computational costs for the policy updates. In addition, Q-learning-based methods such as SAC can introduce model bias during the training process since they estimate gradients using a learned Q-function, whereas policy gradient methods such as PPO have the advantage of directly estimating the gradients of the learning objective. Off-policy algorithms are widely used when data collection is expensive, e.g., when the data can only be obtained from a real robot. The recent literature on RL for humanoid robots relies mainly on simulation data, and hence PPO is widely employed. For other RL settings, such as real-world RL, we envision that more off-policy RL algorithms and novel RL algorithms will play an important role in learning humanoid skills.
VII-A2 Improving Learning Efficiency
@@
Several approaches have been proposed to improve learning efficiency. Curriculum learning expedites training by allowing the policy to achieve simple tasks in the early stage of training and then progressively increasing the task difficulty and complexity [290]. Another approach is to promote exploration. Researchers use curiosity mechanisms, which encourage visiting unexplored states, to intrinsically motivate learning without an explicit reward design [291]. This has been shown to overcome the sparse reward setting and achieve complex loco-manipulation behaviors such as door opening. [281] also incorporates curiosity-based rewards to learn versatile loco-manipulation skills without any motion priors. Lastly, substituting reward terms with constraints in a constrained RL framework can significantly simplify reward tuning while achieving state-of-the-art locomotion performance [292].
VII-A3 Addressing the Sim-to-real Gap
@@
Sim-to-real is another formidable challenge in RL. Nevertheless, RL has been successfully applied to various areas of robotics, notably in quadrupeds [293], where the sim-to-real gap has been consistently overcome. The success of quadrupeds hinges on new investments in infrastructure for affordable hardware and highly parallel physics engines, spearheaded by key players in the robotics field. It is also important to note that quadrupeds benefit from an inherently stable dynamic system similar to manipulators while operating in less complex environments compared to typical loco-manipulation tasks.
@@
In contrast, humanoid loco-manipulation faces steeper sim-to-real challenges. Humanoid robots possess higher DoFs and unstable dynamics, where the center of mass constantly moves out of the support polygon. Therefore, learning whole-body balance control is sensitive to parameters in a physics simulation, underlining the sim-to-real gap due to differences in dynamics between the virtual and real worlds. Additionally, humanoids are expected to perform human-level manipulation tasks where the complex environment and the differences in observation space, such as visual appearances, aggravate the sim-to-real-gap.
@@
To address the sim-to-real challenge, a diverse set of mainstream approaches have been explored for humanoid robots. Domain Randomization (DR) is among the most widely adopted approaches, which varies the properties of a robot model, such as mass, friction, and actuator dynamics, to train a generalized policy robust in the real world. Many humanoid works [282, 42] achieve sim-to-real transfer through DR. While DR is straightforward to set up, policy training is sensitive to the parameter randomization range, inducing laborious tuning: a larger range is challenging for the policy to fit all physical parameters (i.e., fail to learn), and a smaller range does not cover the full spectrum of parameters that can occur in the physical world (i.e., fail to transfer).
@@
System Identification (SI) is another popular approach to enhancing model fidelity by approximating the system’s input-output behavior from real-world data. Real-to-sim techniques use optimization [294] or search [295] to obtain simulation parameters that can best explain the real robot trajectories collected from policy execution. Learning to align simulation behavior with policy rollouts in the real world has also shown success in sim-to-real transfer [296]. However, it is challenging to collect real-world data covering the full space of states and actions, particularly for versatile, safety-critical tasks.
@@
While SI uses real-world data to obtain an accurate model, Domain Adaptation (DA) uses real-world data directly to fine-tune a simulator-trained policy. In DA, the parameter distribution in simulation is defined as the source domain, and parameter distribution in the real world is defined as the target domain. The fine-tuning effectively transfers the policy from the source domain to the target domain. For example, Sim-to-Lab-to-Real [297] develops a two-stage transfer: pre-training in simulation and fine-tuning in the real world. Although only limited hardware data is needed for fine-tuning, safety is still a major concern. Safety filters are often deployed to prevent dangerous movements when collecting real-world data [298].
Despite these efforts to address the sim-to-real gap, a systematic solution remains elusive, as the aforementioned approaches are often case-specific. Against this backdrop, developing physics engines that facilitate real-to-sim construction and sim-to-real transfer is crucial. However, physics engines are constrained by real-time requirements and continue to face challenges in accurately solving contact dynamics. For instance, MuJoCo [214] employs a soft contact model that can lead to unphysical penetration artifacts, while PhysX [215] approximates friction cones using low-fidelity pyramids that widen the sim-to-real gap. Although existing solutions inevitably involve trade-offs between accuracy and efficiency, advancements in contact modeling will significantly narrow the sim-to-real gap.
@@
Key Takeaways: RL provides an effective way to learn novel behaviors for humanoid loco-manipulation. However, in practice, the success of RL often relies on informative representations for both observation and action, extensive reward engineering, curriculum learning design, and a vast amount of trial-and-error experiences to estimate gradients. Consequently, using RL to train robots is almost never practical in the real world, at least for the current stage of development. Therefore, RL policies are predominantly trained in simulations. This makes the sim-to-real gap the Achilles’ heel of RL, significantly dampening its initial promise. Compared with quadruped robots, the sim-to-real gap is particularly challenging for humanoid robots with high DoFs executing complex loco-manipulation tasks. This is why IL, leveraging limited but in-domain real-world data, has gained popularity over pure RL without demonstration data. For further reading in RL, we recommend the survey on learning-based legged locomotion [16] and the practical lessons for training robotic RL agent [299].
TABLE V: Skill Learning Methods Based On Data Source